# JETSCAPE Collaboration

### v1

@@ IV.2 Jet splitting radius
Figure 12: Same as Fig. 8 for $\gamma$-tagged jets from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ ranges. The photon of $160140<p^{\gamma}_{\mathrm{T}}<$ 160 GeV is triggered with isolation requirement, relative azimuth angle cut.
Figure 12 shows our prediction for the medium modification of the $r_{g}$ distribution for $\gamma$-tagged jets in central Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ cuts. Unlike the $z_{g}$ distribution, the $r_{g}$ distribution exhibits a strong $x_{J\gamma}$ dependence. For the case with large $x_{J\gamma}$ cuts, the selection bias effect due to the $r_{g}$-dependence of energy loss appears prominently, resulting in a strong suppression at large $r_{g}$ values. Here, it should be noted that the effect of the strong $p^{\mathrm{jet}}_{T}$ dependence on the $r_{g}$ distribution of jets in vacuum is diminished, since, by setting an uppercut on $p^{\gamma}_{T}$, the feeddown from jets originating from parent partons with larger initial $p_{T}$ is suppressed.
The selection bias effect is counterbalanced by the strong broadening effects on hard splittings, attributable to the large fraction of quark jets, in the small-to-mid $r_{g}$ range. When the value of the lower $x_{J\gamma}$ cut is reduced to weaken the selection bias effect, the large $r_{g}$ suppression diminishes, and concurrently, even a bump structure due to broadening emerges at mid $r_{g}$ ($\lessapprox 0.1$). This manifestation of the broadening effect in the $r_{g}$ distribution represents a distinct characteristic of $\gamma$-tagged jets, contracting with the behavior of inclusive jets, where only a monotonous pattern entirely dominated by selection bias can be seen. Particularly, by varying the $x_{J\gamma}$ cut, one can control the selection bias effect and study the broadening effect more quantitatively. Therefore, $\gamma$-tagged jets provide a modification pattern in the $r_{g}$ distribution that is more suitable for investigating structural modifications on the hard partonic splittings.
@@ IV.3 Relative transverse momentum of jet splittings
Figure 13: Distributions of relative transverse momentum of jet splittings $k_{T,g}$ normalized by the number of all triggered jets for the leading jets in events generated with pgun. The jet shower evolution is performed by vacuum matter for the parent parton having a $E_{\mathrm{init}}=140$ GeV. Jets are reconstructed with $R=0.4$ at midrapidity $|\eta_{\mathrm{jet}}|<2.0$. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The Soft Drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
@@
Figure 14: Ratios of $k_{T,g}$ distributions for the leading jets in events with the parent parton having a fixed initial energy $E_{\mathrm{init}}=140$ GeV generated by pgun. The jet shower evolution in the QGP medium produced in 0%–10% Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is performed by matter$+$lbt. All the setups are the same as in Fig. 2.
The medium modifications in the $k_{T,g}$ distribution for the pgun jets are shown in Fig. 14. In the results for both quark and gluon jets, suppression at large $k_{T,g}$ values is seen. As seen in Eq. (10), $k_{T,g}$ increases with increasing $r_{g}$. Thus, one might naively think that this suppression is attributed to the same selection bias effect as in the $r_{g}$ distribution: larger energy loss in jets with larger $k_{T,g}$. However, this is not only the cause, as evidenced by the persistence of the suppression at large-$k_{T,g}$ even when the $p^{\mathrm{jet}}_{T}$ cut is lowered to reduce the effect of hard splitting angle ($r_{g}$) dependence on energy loss. This suppression at large $k_{T,g}$ is actually brought also by the loss of $p_{T,2}$: large-angle soft radiations in the later stage can be trimmed by either the jet cone or the Soft Drop grooming.
The $p_{T,2}$ loss effect becomes apparent when focusing on the gluon jet with the smallest presented $p^{\mathrm{jet}}_{T}$ cut ($p^{\mathrm{jet}}_{T}/E_{\mathrm{init}}>0.2$), taking into account the results of the $r_{g}$ modification shown in Fig. 8. In this instance, whether due to energy loss effects or broadening, there is almost no modification observed in the $r_{g}$ distribution of gluon jets. Thus, from Eq. (10), the modification in the $k_{T,g}$ must be attributed to the change in $p_{T,2}$, in particular, for the case with the small $p^{\mathrm{jet}}_{T}$ cut.
In the small to mid $k_{T,g}$ ($\lessapprox 2$ GeV) region, similar to what was observed in the modification of the $r_{g}$ distribution, a bump structure due to broadening is observed for quark jets. For the gluon jet case, the modification pattern is dominated by the effect of $p_{T,2}$ loss, with the $r_{g}$-dependence of jet energy loss also exerting a slight influence in cases of large $p^{\mathrm{jet}}_{T}$ cuts.
@@
Figure 16: Same as Fig. 13 for jets from hard scatterings in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, isolation requirement, relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
In Fig. 16, the comparison between the $k_{T,g}$ distributions for $\gamma$-tagged jets and inclusive jets in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV is shown for the same $p^{\mathrm{jet}}_{T}$ cuts. Inclusive jets exhibit a broader transverse momentum in hard splittings compared to $\gamma$-tagged jets, attributed to a larger gluon jet fraction, but this difference diminishes as $p^{\mathrm{jet}}_{T}$ increases.
@@
Figure 17: Same as Fig. 14 for jets from hard scatterings at $\sqrt{s_{NN}}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, isolation requirement, relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
Figure 17 presents the medium modification of the $k_{T,g}$ distribution for $\gamma$-tagged jets and inclusive jets. For both $\gamma$-tagged jets and inclusive jets, the overall modification exhibits a narrowing pattern mostly dominated by the $p_{T,2}$ loss of the prongs. Although no significant bump structure is observed, in the case of $\gamma$-tagged jets, a subtle indication of momentum-broadening effects on quark jets in the lbt phase can be discerned as a flattening on the small $k_{T,g}$ side.
@@
Figure 18: Same as Fig. 14 for $\gamma$-tagged jets from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ ranges. The photon of $160140<p^{\gamma}_{\mathrm{T}}<$ 160 GeV is triggered with isolation requirement, relative azimuth angle cut.
@@
Finally, the prediction for the $x_{J\gamma}$-dependent medium modification of the $k_{T,g}$ distribution for $\gamma$-tagged jets in central Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is shown in Fig. 18. Given that the suppression at large $k_{T,g}$ persists even with a reduced $x_{J\gamma}$ cut ($x_{J\gamma}>0.2$), it is evident that jets with large $k_{T,g}$ cannot recover, even when jets with significant energy loss are included, due to the presence of $p_{T,2}$ loss. On the small $k_{T,g}$ side, a very flat pattern is seen as a consequence of the competition between the transverse momentum broadening of quark jets in lbt and the other effects narrowing the distribution.
@@ IV.4 Groomed jet mass
Figure 19: Distributions of the relative transverse momentum of jet splittings $m_{g}$ normalized by the number of all triggered jets for the leading jets in events generated with pgun. The jet shower evolution is performed by vacuum matter for the parent parton having a $E_{\mathrm{init}}=140$ GeV. Jets are reconstructed with $R=0.4$ at midrapidity $|\eta_{\mathrm{jet}}|<2.0$. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The Soft Drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
In Fig. 19, the $m_{g}$ distributions for the vacuum pgun jets are shown. Here, again, due to the larger virtuality at the initial hard scattering production, gluon jets have a broader distribution with larger $m_{g}$ than quark jets. For the vacuum case with fixed $E_{\mathrm{init}}=140$ GeV, no significant $p^{\mathrm{jet}}_{T}$ cut dependence is seen for the jet cone size $R=0.4$.
Figure 20: Ratios of $m_{g}$ distributions for the leading jets in events with the parent parton having a fixed initial energy $E_{\mathrm{init}}=140$ GeV generated by pgun. The jet shower evolution in the QGP medium produced in 0%–10% Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is performed by matter$+$lbt. All the setups are the same as in Fig. 2.
The medium modification of $m_{g}$ distribution for jets from the pgun simulations with fixed $E_{\mathrm{init}}$ is shown in Fig. 20. In the case of quark jets, a clear bump structure is observed due to the shift of small $m_{g}$ to mid $m_{g}$ caused by medium effects. On the other hand, such behavior is not observed for gluon jets, which exhibit a monotonic trend. Besides, for both quark and gluon jets with large $p^{\mathrm{jet}}_{T}$ cuts, suppression at large $m_{g}$ is observed, attributed to the greater energy loss in jets with larger masses, which results in wider splittings. Then, with the small $p^{\mathrm{jet}}_{T}$ cuts, the suppression becomes slightly moderate at mid $m_{g}$ (10absent10$\approx$ 10≈ 1015151515 GeV), yet still dominates the modification. This is because, analogous to the case for $k_{T,g}$, the jet energy loss—primarily due to large-angle energy emission outside of the jet cone—results in the loss of $m_{g}$.
@@
Figure 22: Same as Fig. 19 for jets from hard scatterings in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, isolation requirement, relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
The $m_{g}$ distributions for $\gamma$-tagged jets and inclusive jets in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV are compared in Fig. 22. The inclusive jet results exhibit a broader distribution in the $m_{g}$ due to the larger gluon jet fraction, but the difference diminishes as $p^{\mathrm{jet}}_{T}$ increases, similarly to the $r_{g}$ and $k_{T,g}$ distributions.
@@
Figure 23: Same as Fig. 20 for jets from hard scatterings at $\sqrt{s_{NN}}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for the inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) and $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, isolation requirement, relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
The medium modification is shown in Fig. 23. For inclusive jets, the modification pattern is dominated by the shift towards smaller $m_{g}$, attributed to larger energy loss with wider angle profiles, and does not show any clear manifestation of mass gain for the hard splittings. In the case of $\gamma$-tagged jets, a bump due to the mass gain for quark jets in the lbt phase appears on the small $m_{g}$ side, while a similar suppression to that of inclusive jets is observed at large $m_{g}$.
@@
Figure 24: Same as Fig. 20 for $\gamma$-tagged jets from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ ranges. The photon of $160140<p^{\gamma}_{\mathrm{T}}<$ 160 GeV is triggered with isolation requirement, relative azimuth angle cut.
@@
We present our prediction for the $x_{J\gamma}$-dependent medium modification of the $m_{g}$ distribution for $\gamma$-tagged jets in Fig. 24. A bump structure associated with mass gain from the medium is observed in hard splittings. At large $m_{g}$, jets with larger masses experience greater energy loss, leading to wider splittings and suppression at large $x_{J\gamma}$ cuts. Reducing the $x_{J\gamma}$ cut includes contributions from jets with significant energy loss, weakening the suppression, but it persists even at the smallest $x_{J\gamma}$ cut ($x_{J\gamma}>0.2$). This is attributed to the loss of $m_{g}$ caused by energy emitted at large angles outside the jet cone or soft components groomed away by the Soft Drop procedure.

### v2

@@ IV.2 Jet splitting radius
Figure 14: Same as Fig. 9 for $\gamma$-tagged jets from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ ranges. The photon of $140<p^{\gamma}_{\mathrm{T}}<160$ GeV is triggered with the isolation requirement, and the relative azimuth angle cut.
Figure 14 shows our prediction for the medium modification of the $r_{g}$ distribution for $\gamma$-tagged jets in central Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ cuts. Unlike the $z_{g}$ distribution, the $r_{g}$ distribution exhibits a strong $x_{J\gamma}$ dependence. For the case with large $x_{J\gamma}$ cuts, the selection bias effect due to the $r_{g}$ dependence of energy loss appears prominently, resulting in a strong suppression at large $r_{g}$ values. Here, it should be noted that the effect of the strong $p^{\mathrm{jet}}_{T}$ dependence on the $r_{g}$ distribution of jets in vacuum is diminished, since, by setting an uppercut on $p^{\gamma}_{T}$, the feeddown from jets originating from parent partons with larger initial $p_{T}$ is suppressed.
The selection bias effect is counterbalanced by the strong broadening effects, attributable to the large fraction of quark jets, in the small-to-mid $r_{g}$ range. When the value of the lower $x_{J\gamma}$ cut is reduced to weaken the selection bias effect, the large $r_{g}$ suppression diminishes, and concurrently, even a bump structure due to broadening emerges at mid $r_{g}$ ($\lessapprox 0.1$). This manifestation of the broadening effect in the $r_{g}$ distribution represents a distinct characteristic of $\gamma$-tagged jets, in contrast with the behavior of inclusive jets, where only a monotonic pattern entirely dominated by selection bias can be seen. Particularly, by varying the $x_{J\gamma}$ cut, one can control the selection bias effect and study the broadening effect more quantitatively. Therefore, $\gamma$-tagged jets provide a modification pattern in the $r_{g}$ distribution that is more suitable for investigating the structural modifications on the soft-drop-reconstructed hard branchings.
@@ IV.3 Relative transverse momentum of jet splittings
Figure 15: Distributions of relative transverse momentum of jet splittings $k_{T,g}$ normalized by the number of all triggered jets for the leading jets in events generated with pgun. The jet shower evolution is performed by vacuum matter for the parent parton having $E_{\mathrm{init}}=140$ GeV. Jets are reconstructed with $R=0.4$ at midrapidity $|\eta_{\mathrm{jet}}|<2.0$. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The soft-drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
@@
Figure 16: Ratios of $k_{T,g}$ distributions for the leading jets in events with the parent parton having a fixed initial energy $E_{\mathrm{init}}=140$ GeV generated by pgun. The jet shower evolution in the QGP medium produced in 0%–10% Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is performed by matter++lbt. All the setups are the same as in Fig. 2.
The medium modifications in the $k_{T,g}$ distribution for the pgun jets are shown in Fig. 16. In the results for both quark and gluon jets, suppression at large $k_{T,g}$ values is seen. As seen in Eq. (10), $k_{T,g}$ increases with increasing $r_{g}$. Thus, one might naively think that this suppression is attributed to the same selection bias effect as in the $r_{g}$ distribution: larger energy loss in jets with larger $k_{T,g}$. However, this is not only the cause, as evinced by the persistence of the suppression at large $k_{T,g}$ even when the $p^{\mathrm{jet}}_{T}$ cut is lowered to reduce the effect of hard splitting angle ($r_{g}$) dependence on energy loss. This suppression at large $k_{T,g}$ is actually brought also by the loss of $p_{T,2}$: large-angle soft radiations in the later stage can be trimmed by either the jet cone or the soft-drop grooming.
The $p_{T,2}$ loss effect becomes apparent when focusing on the gluon jet with the smallest presented $p^{\mathrm{jet}}_{T}$ cut ($p^{\mathrm{jet}}_{T}/E_{\mathrm{init}}>0.2$), taking into account the results of the $r_{g}$ modification shown in Fig. 9. In this instance, whether due to energy-loss effects or broadening, there is almost no modification observed in the $r_{g}$ distribution of gluon jets. Thus, from Eq. (10), the modification in the $k_{T,g}$ must be attributed to the change in $p_{T,2}$, in particular, for the case with the small $p^{\mathrm{jet}}_{T}$ cut.
In the small to mid $k_{T,g}$ ($\lessapprox 2$ GeV) region, similar to what was observed in the modification of the $r_{g}$ distribution, a bump structure indicating broader hard branchings is observed for quark jets. For the gluon jet case, the modification pattern is dominated by the effect of $p_{T,2}$ loss, with the $r_{g}$ dependence of jet energy loss also exerting a slight influence in cases of large $p^{\mathrm{jet}}_{T}$ cuts.
Figure 17: Same as Fig. 16 for matter++lbt simulations in which the medium effects from recoils and holes are artificially removed.
Figure 17 shows the matter++lbt results without the contributions from recoils and holes. For gluon jets, the results remain almost unchanged compared with those with the full medium-response contribution. In contrast, for quark jets, the bump observed in the full medium-response results disappears, and their qualitative behavior becomes identical to that of gluon jets, exhibiting a monotonic decrease as a function of $k_{T}$. Furthermore, unlike in the case of $r_{g}$, no residual bump is visible, indicating that the bump in $k_{T,g}$ originates entirely from recoils, while the bump in $r_{g}$ observed even without recoils is caused by soft branchings with small $k_{T,g}$.
@@
Figure 19: Same as Fig. 15 for jets from hard scatterings in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, the isolation requirement, the relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
In Fig. 19, the comparison between the $k_{T,g}$ distributions for $\gamma$-tagged jets and inclusive jets in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV is shown for the same $p^{\mathrm{jet}}_{T}$ cuts. Inclusive jets exhibit a broader transverse momentum in hard splittings compared with $\gamma$-tagged jets, attributed to a larger gluon jet fraction, but this difference diminishes as $p^{\mathrm{jet}}_{T}$ increases.
@@
Figure 20: Same as Fig. 16 for jets from hard scatterings at $\sqrt{s_{NN}}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, the isolation requirement, the relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
Figure 20 presents the medium modification of the $k_{T,g}$ distribution for $\gamma$-tagged jets and inclusive jets. For both $\gamma$-tagged jets and inclusive jets, the overall modification exhibits a narrowing pattern mostly dominated by the $p_{T,2}$ loss of the prongs. Although no pronounced bump structure is observed, $\gamma$-tagged jets exhibit a subtle indication of broadening in the hard branchings of quark jets due to recoils generated in the lbt phase, manifesting as a slight flattening on the small-$k_{T,g}$ side.
@@
Figure 21: Same as Fig. 16 for $\gamma$-tagged jets from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ ranges. The photon of $140<p^{\gamma}_{\mathrm{T}}<160$ GeV is triggered with the isolation requirement, and the relative azimuth angle cut.
@@
Finally, the prediction for the $x_{J\gamma}$-dependent medium modification of the $k_{T,g}$ distribution for $\gamma$-tagged jets in central Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is shown in Fig. 21. Given that the suppression at large $k_{T,g}$ persists even with a reduced $x_{J\gamma}$ cut ($x_{J\gamma}>0.2$), it is evident that jets with large $k_{T,g}$ cannot recover, even when jets with significant energy loss are included, due to the presence of $p_{T,2}$ loss. On the small-$k_{T,g}$ side, a very flat pattern appears as a result of the competition between broad branchings generated by recoils in quark jets during the lbt phase and other effects that narrow the distribution.
@@ IV.4 Groomed jet mass
Figure 22: Distributions of the relative transverse momentum of jet splittings $m_{g}$ normalized by the number of all triggered jets for the leading jets in events generated with pgun. The jet shower evolution is performed by vacuum matter for the parent parton having $E_{\mathrm{init}}=140$ GeV. Jets are reconstructed with $R=0.4$ at midrapidity $|\eta_{\mathrm{jet}}|<2.0$. The results are shown for quark jets (solid) and gluon jets (dashed) with different $p^{\mathrm{jet}}_{T}$ triggers, 112, 84, 56, and 28 GeV. The soft-drop parameters are $z_{\mathrm{cut}}=0.2$ and $\beta=0$.
In Fig. 22, the $m_{g}$ distributions for the vacuum pgun jets are shown. Here, again, due to the larger virtuality at the initial hard scattering production, gluon jets have a broader distribution with larger $m_{g}$ than quark jets. For the vacuum case with fixed $E_{\mathrm{init}}=140$ GeV, no significant $p^{\mathrm{jet}}_{T}$-cut dependence is seen for the jet cone size $R=0.4$.
Figure 23: Ratios of $m_{g}$ distributions for the leading jets in events with the parent parton having a fixed initial energy $E_{\mathrm{init}}=140$ GeV generated by pgun. The jet shower evolution in the QGP medium produced in 0%–10% Pb-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV is performed by matter++lbt. All the setups are the same as in Fig. 2.
The medium modification of $m_{g}$ distribution for jets from the pgun simulations with fixed $E_{\mathrm{init}}$ is shown in Fig. 23. In the case of quark jets, a clear bump structure is observed due to the shift of small $m_{g}$ to mid $m_{g}$ caused by medium effects. On the other hand, such behavior is not observed for gluon jets, which exhibit a monotonic trend. Besides, for both quark and gluon jets with large $p^{\mathrm{jet}}_{T}$ cuts, suppression at large $m_{g}$ is observed, attributed to the greater energy loss in jets with larger masses, which results in wider splittings. Then, with the small $p^{\mathrm{jet}}_{T}$ cuts, the suppression becomes slightly moderate at mid $m_{g}$ ($\approx 10$1515 GeV), yet still dominates the modification. This is because, analogous to the case for $k_{T,g}$, the jet energy loss—primarily due to large-angle energy emission outside of the jet cone—results in the loss of $m_{g}$ [97].
Figure 24: Same as Fig. 23 for matter++lbt simulations in which the medium effects from recoils and holes are artificially removed.
Figure 24 shows the results with the contributions from recoils and holes artificially removed. As in the case of $k_{T,g}$, the results for gluon jets remain almost unchanged compared with those with the full medium response. For quark jets, however, the bump observed with the full medium response disappears, and their qualitative behavior becomes identical to that of gluon jets, exhibiting a monotonic decrease as a function of $m_{g}$. The slight recoil-independent broadening seen in $r_{g}$ originates from soft branchings that do not produce a bump in $m_{g}$.
@@
Figure 26: Same as Fig. 22 for jets from hard scatterings in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) and inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, the isolation requirement, the relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
The $m_{g}$ distributions for $\gamma$-tagged jets and inclusive jets in $p$-$p$ collisions at $\sqrt{s}=5.02$ TeV are compared in Fig. 26. The inclusive jet results exhibit a broader distribution in $m_{g}$ due to the larger gluon jet fraction, but the difference diminishes as $p^{\mathrm{jet}}_{T}$ increases, similarly to the $r_{g}$ and $k_{T,g}$ distributions.
@@
Figure 27: Same as Fig. 23 for jets from hard scatterings at $\sqrt{s_{NN}}=5.02$ TeV, generated by pythiagun with ISR and MPI. The results are shown for the inclusive jets (dashed) from inclusive hard processes (HardQCD:all=on+PromptPhoton:all=on) and $\gamma$-tagged jets (solid) from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 with different $p^{\mathrm{jet}}_{\mathrm{T}}$ triggers. For $\gamma$-tagged jets, the isolation requirement, the relative azimuth angle cut, and the additional cut of $p^{\mathrm{jet}}_{\mathrm{T}}<p^{\gamma}_{\mathrm{T}}$ are imposed.
@@
The medium modification is shown in Fig. 27. For inclusive jets, the modification pattern is dominated by the shift toward smaller $m_{g}$, attributed to larger energy loss with wider angle profiles. In the case of $\gamma$-tagged jets, a bump due to the mass gain for quark jets via recoils in the lbt phase appears on the small $m_{g}$ side, while a similar suppression to that of inclusive jets is observed at large $m_{g}$.
@@
Figure 28: Same as Fig. 23 for $\gamma$-tagged jets from prompt photon-generating hard processes (HardQCD:all=off+PromptPhoton:all=on) generated at leading order by pythia 8 at $\sqrt{s_{NN}}=5.02$ TeV for different $x_{J\gamma}$ ranges. The photon of $140<p^{\gamma}_{\mathrm{T}}<160$ GeV is triggered with the isolation requirement, and the relative azimuth angle cut.
@@
We present our prediction for the $x_{J\gamma}$-dependent medium modification of the $m_{g}$ distribution for $\gamma$-tagged jets in Fig. 28. A bump structure associated with mass gain via recoils from the medium at small virtuality is observed in soft-drop-reconstructed hard branchings. At large $m_{g}$, jets with larger masses experience greater energy loss, leading to wider splittings and suppression at large $x_{J\gamma}$ cuts. Reducing the $x_{J\gamma}$ cut includes contributions from jets with significant energy loss, weakening the suppression, but it persists even at the smallest $x_{J\gamma}$ cut ($x_{J\gamma}>0.2$). This is attributed to the loss of $m_{g}$ caused by energy emitted at large angles outside the jet cone or soft components groomed away by the soft-drop procedure.